# The Nonstationary Newsvendor with (and without) Predictions

### v3

@@ Appendix H Unknown $v$ and Known $a$
             $\delta_{t}=(C_{\max}-(b_{t}(d_{t}-q^{\mathrm{shrinking}}_{t})^{+}+h_{t}(q^{\mathrm{shrinking}}_{t}-d_{t})^{+}))/p^{\mathrm{shrinking}}(t)$;
             $w^{\mathrm{shrinking}}(t+1)=w^{\mathrm{shrinking}}(t)\exp{(\gamma\delta_{t}/2)}$;
@@
            $\delta_{t}=(C_{\max}-(b_{t}(d_{t}-q^{\mathrm{prediction}}_{t})^{+}+h_{t}(q^{\mathrm{prediction}}_{t}-d_{t})^{+}))/p^{\mathrm{prediction}}(t)$;
             $w^{\mathrm{prediction}}(t+1)=w^{\mathrm{prediction}}(t)\exp{(\gamma\delta_{t}/2)}$.
@@ Appendix I Experiment Details
In our experiment, we first generated a demand sequence for 30 time periods where each time period is drawn uniformly at random between 80 and 120. This was treated as the historical observations and was fixed throughout the experiment. Then for each set of parameters $(\alpha,\beta,\gamma,L)$, which we will specify later in each case, we used triple exponential smoothing to generate the mean of demands for 365 time periods, where at each time period we also added a random Gaussian noise with mean equals to 0 and variance equals to 5. Finally the true demand at each time period was generated as a Poisson variable with the corresponding mean.

### v4

@@ Appendix H Unknown $v$ and Known $a$
             $\delta_{t}=(b_{t}(d_{t}-q^{\mathrm{shrinking}}_{t})^{+}+h_{t}(q^{\mathrm{shrinking}}_{t}-d_{t})^{+})/p^{\mathrm{shrinking}}(t)$;
             $w^{\mathrm{shrinking}}(t+1)=w^{\mathrm{shrinking}}(t)\exp{(-\gamma\delta_{t}/2)}$;
@@
            $\delta_{t}=(b_{t}(d_{t}-q^{\mathrm{prediction}}_{t})^{+}+h_{t}(q^{\mathrm{prediction}}_{t}-d_{t})^{+})/p^{\mathrm{prediction}}(t)$;
             $w^{\mathrm{prediction}}(t+1)=w^{\mathrm{prediction}}(t)\exp{(-\gamma\delta_{t}/2)}$.
@@ Appendix I Experiment Details
In our experiment, we first generated a demand sequence for 30 time periods where the demand at each time period is drawn uniformly between 80 and 120. This was treated as the historical observations and was fixed throughout the experiment. Then for each set of parameters $(\alpha,\beta,\gamma,L)$, which we will specify later in each case, we used triple exponential smoothing to generate the mean of demands for 365 time periods, where at each time period we also added a random Gaussian noise with mean equals to 0 and variance equals to 5. Finally the true demand at each time period was generated as a Poisson variable with the corresponding mean.
@@
Appendix J Additional Experimental Results
In our experiments on real data, we used four popular forecasting method to generate predictions for each instance: Exponential Smoothing (Holt Winters), ARIMA, Prophet, and LightGBM. Experiments on the datasets yielded the histograms in Fig. 4. To show the performance of PERP is robust to the forecasting method, in Fig. 6 we further divide the three histograms in Fig. 4 into twelve histograms separated by the four forecasting methods.
(a) GAPs with Exponential Smoothing forecasts. Left to right: Rossmann, Wikipedia, Restaurant.
(b) GAPs with ARIMA forecasts. Left to right: Rossmann, Wikipedia, Restaurant.
(c) GAPs with Prophet forecasts. Left to right: Rossmann, Wikipedia, Restaurant.
(d) GAPs with ELightGBM forecasts. Left to right: Rossmann, Wikipedia, Restaurant.
Figure 6. Histograms of GAPs divided by forecasting methods across the Rossmann dataset, the Wikipedia dataset, and the Restaurant dataset.